Teal, Inc. began work on a $6,550,000 contract in 2020 to construct an office building. During 2020, Teal, Inc. incurred costs of $1,817,440, billed its customers for $1,316,000, and collected $1,039,000. At December 31, 2020, the estimated additional costs to complete the project total $3,094,560.

Prepare Teal’s 2020 journal entries using the percentage-of-completion method.
Answer #1

Answer

  • Requirement asked

Date

Accounts title

Debit

Credit

2020

Construction in Progress

$1,817,440

   Various accounts

$1,817,440

(construction cost incurred)

2020

Accounts receivables

$1,316,000

Billing on Construction contracts

$1,316,000

(amount billed during the year)

2020

Cash

$1,039,000

Accounts receivables

$1,039,000

(cash received)

2020

Construction in Progress

$606,060

Cost of construction

$1,817,440

Revenue from Long Term contract

$2,423,500

(revenue recognised)

--Working

Working

2020

A

Contract Price

$6,550,000

B

Cost Incurred to Date

$1,817,440

C

Estimated cost yet to be incurred to complete the contract

$3,094,560

D = B+C

Total Cost

$4,912,000

E = (B/D) x 100

% of Completion

37.00%

F = A x E

Revenue to date

$2,423,500

G

Revenue of Previous year

$0

H = F - G

Net Revenue this year

$2,423,500

I (=B)

Cost to date

$1,817,440

J

Cost to date of previous year

$0

K = I - J

Net Cost for the year

$1,817,440

L = H - K

Gross Profits

$606,060